OpenStudy (anonymous):

Simplify quantity 6 times x plus 24 all over 6 PLEASE HELP!!!!!!!!!!! DO I MULTIPLY SIX ON BOTH SIDES?

OpenStudy (lalaly):

both sides? u dont have any sides, its only one fraction

OpenStudy (anonymous):

so what should i do? @lalaly

OpenStudy (lalaly):

\[\frac{6x+24}{6}\] this is equivalent to\[\frac{6x}{6}+\frac{24}{6}\]now whats 6x/6?

OpenStudy (goformit100):

Great Work madam

OpenStudy (lalaly):

:) thanks @goformit100 :D

OpenStudy (anonymous):

Would the answer be x+4? @lalaly

OpenStudy (lalaly):

yep

OpenStudy (anonymous):

Thank you!! can you help me with one more please? @lalaly

OpenStudy (lalaly):

sure:)

OpenStudy (anonymous):

Okay this one is a little different. Solve for y: 2 over 3y + 4 = 6 @lalaly

OpenStudy (lalaly):

or u want me to show u?

OpenStudy (anonymous):

multiply by the Reciprocal? @lalaly

OpenStudy (lalaly):

yeah exactly can u do that?

OpenStudy (anonymous):

yes \[2\div3 \times 2\div3 =\frac{ 4 }{ 6 }\] now what?

OpenStudy (anonymous):

@lalaly

OpenStudy (anonymous):

subtract 4 from both sides? @lalaly

OpenStudy (lalaly):